# LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation

### v1

@@ Abstract
The global COVID-19 pandemic underscored major deficiencies in traditional healthcare systems, hastening the advancement of online medical services, especially in medical triage and consultation. However, existing studies face two main challenges. First, the scarcity of large-scale, publicly available, domain-specific medical datasets due to privacy concerns, with current datasets being small and limited to a few diseases, limiting the effectiveness of triage methods based on Pre-trained Language Models (PLMs). Second, existing methods lack medical knowledge and struggle to accurately understand professional terms and expressions in patient-doctor consultations. To overcome these obstacles, we construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC), comprising a Coarse-grained Triage dataset with 439,630 samples, a Fine-grained Diagnosis dataset with 199,600 samples, and a Medical Consultation dataset with 472,418 items, thereby addressing the data shortage in this field. Moreover, we further propose a novel triage system that combines BERT-based supervised learning with prompt learning, as well as a GPT-based medical consultation model using reinforcement learning. To enhance domain knowledge acquisition, we pre-trained PLMs using our self-constructed background corpus. Experimental results on the LCMDC demonstrate the efficacy of our proposed systems. Our corpora and codes are available at: Corpora and Codes
LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation
Xinyuan Wang  and Haozhou Li  and Dingfang Zheng  and Qinke Peng Xi’an Jiaotong University {wxy0713, lihaozhou1126, zdf0921}@stu.xjtu.edu.cn qkpeng@mail.xjtu.edu.cn
1 Introduction
@@
The outbreak and spread of COVID-19 have placed immense pressure on global healthcare systems, exposing numerous shortcomings in traditional healthcare services Sun et al. (2021). In response to the urgent need for pandemic control, online medical consultations and tele-medicine have become crucial for maintaining healthcare services Jnr (2020). Between July and December 2022, 58.5% of U.S. adults used the Internet to seek health or medical information, 41.5% communicated with a doctor or doctor’s office, and 46.1% checked medical test resultsWang and Cohen (2023). According to the China Internet Network Information Center (CNNIC), as of December 2023, the number of Internet medical users in China reached 414 million. This trend reflects the growing reliance on digital platforms for healthcare.
@@
The rise of online healthcare platforms has made medical triage and intelligent Q&A key research areas, helping reduce the burden on healthcare systems by providing remote medical advice and reducing unnecessary hospital visits. These technologies enhance resource efficiency, minimize cross-infection risks, and improve access to healthcare services, especially in remote areas.
@@
However, most existing studies face two main challenges. Firstly, due to data privacy concerns, obtaining large-scale public medical triage and consultation datasets is difficult. Although some studies have collected corpora for automatic diagnosis Wei et al. (2018); Lin et al. (2019); Xu et al. (2019) and dialogue generation Liu et al. (2022), these datasets are limited to a few specific diseases and are small in size Zeng et al. (2020), making them inadequate for training intelligent medical diagnosis models based on pre-trained language models. Secondly, patient-doctor consultations usually utilize distinct expressions and professional terms, making it difficult for existing methods that lack domain-specific knowledge to comprehend medically pertinent information Li et al. (2024).
@@
To address the limitations of existing datasets, we have constructed the Large-scale Chinese Medical Dialogue Corpora (LCMDC) by extracting patient-doctor conversations from the "Quick Doctor" website  [^1] https://www.120ask.com/. Our work addresses two key needs in online healthcare: task-oriented medical triage and open-ended consultation systems. We developed a coarse-grained triage dataset with over 430,000 patient-doctor dialogues across 14 departments and a fine-grained diagnosis dataset with 200,000 entries covering 120 diseases. For medical consultation, we compiled over 470,000 question-answer pairs. Additionally, we collected disease information datasets to enrich Pre-trained Language Models (PLMs) with medical knowledge for pre-training.
@@
Nowadays, natural language processing techniques are increasingly effective in intelligent medical triage and dialogue generation tasksSun et al. (2017). Specifically, neural network-based text classification algorithms with attention mechanisms have been successfully applied to medical triage and diagnosisWei et al. (2018); Lin et al. (2019); Xu et al. (2019). However, existing studies often concentrate on a limited number of diseases and struggle with imbalanced label distributions in many-class medical diagnoses Zeng et al. (2020). Although some pre-trained language models, such as BERT and ERNIE, have been employed in automatic triage systems Wang et al. (2021); Wang and Feng (2022), they lack pre-training on medical data, hindering their ability to grasp domain-specific knowledge Li et al. (2024). To overcome these limitations, we propose a novel triage system that combines PLM-based supervised learning with prompt learning methods. We initially pre-train BERT using a self-built medical background corpus, then employ medical BERT, BiLSTM, and Dendrite Network Liu and Wang (2021) to represent the input questions for classification in our Triage datasets. Additionally, we incorporate prompt learning to mitigate the many-label imbalance issue, enhancing prediction accuracy for rare diseases.
@@
In the medical text generation task, sequence-to-sequence models, including LSTM-based Seq2Seq methods Liu et al. (2022) and transformer-based PLMs like BERT-GPT Zeng et al. (2020), have been widely used, and achieved promising results. However, traditional LSTM models, which generate output solely based on input sequences, struggle with insufficient information utilization due to their gating mechanisms. Similarly, conventional PLMs often fail to effectively comprehend medical information within patient-doctor conversations. Additionally, these methods typically lack supplementary information for enhancement, resulting in poor performance in generating patient-doctor dialogues. To address these challenges, we re-train GPT-2 using our collected medical background corpus to internalize domain-specific knowledge. We also construct a domain knowledge graph and extract relevant content about each user’s question as supplementary input information to enhance the model. Finally, we train our model using our Medical Consultation Dataset and fine-tune it with a reinforcement learning method to simulate human doctors in providing responses.
@@
  • Developed a dialogue model construction framework including knowledge incorporation and input supplementation, improving performance across multiple evaluation metrics.
2 Chinese Medical Corpus
@@
2.1 Data Collection
To make the model more adaptable to conversational and ambiguous medical descriptions, we claw the original data from the online medical consultation platform named "Quick Doctor". Reasons are shown in Appendix B.1.
@@ 2.1 Data Collection
We collect millions of entries containing doctor responses and conduct statistical analysis and visualization based on different labels to better understand the original data. Our analysis, illustrated in Figure 1, explores the gender and age of users. The results show that female users significantly outnumber male users. Predominantly, young individuals aged 10-40, who are more likely to incorporate the Internet into their daily routines, made the inquiries.
2.2 Intelligent Triage Dataset
@@
We organize the original corpus into two datasets after data preprocessing: 1) a Coarse-grained Triage dataset with primary department labels for higher accuracy in quickly identifying potential conditions and directing patients to the appropriate department, and 2) a Fine-grained Diagnosis dataset with tertiary department labels for more precise guidance, thereby helping users access detailed medical knowledge and improving query efficiency. The Coarse-grained dataset has over 430,000 entries with 14 categories, while the Fine-grained dataset has over 190,000 entries with 120 categories, as shown in Table 1. Based on these datasets, we further propose an intelligent triage algorithm based on supervised learning and prompt-based learning and then evaluated its effectiveness.
Table 1: Statistic of Intelligent Triage Datasets
@@
2.3 Medical Consultation Dataset
@@
Table 2: Statistic of Medical Consulting Dataset
@@
Specifically, each sample in this dataset comprises a patient consultation and a doctor’s response, serving respectively as the model’s input and prediction target. We conducted data cleaning to remove entries that lacked either questions or answers and eliminated entries where the questions and responses were shorter than 10 characters, thereby enriching the training data with more substantive content. After data preprocessing, the Medical Consultation Dataset contains a total of 472,418 samples, with an average question length of 57.86 characters and an average response length of 92.57 characters. The statistical results of the dataset are presented in Table 2.
3 Intelligent Triage System
@@ 3.1 Problem Definition
where $p\left\{{c_{1},~{}c_{2},\ldots,c_{k_{l}}}\right\}$ means the probability on each label.
3.2 Supervised Learning Triage System
@@
3.3 Prompt Learning Triage System
@@
4 Medical Consultation System
@@ 4.1 Problem Definition
4.2 Overview
Based on the GPT-2 model, we re-train the model using the collected medical encyclopedia text data to internalize knowledge, making it more suitable for tasks in the medical domain. Next, we construct a medical knowledge graph and use a search-and-match approach to find relevant knowledge to each question, providing external information as supplementary input text. Then, we train the model using dialogue data and finally fine-tune it using reinforcement learning.
@@ 4.2 Overview
4.3 Knowledge Injection Answer Generation
@@
$L\left(Sentence\right)=\sum\limits_{n=K}^{N}p\left({token}_{n}\,\middle|\,{token}_{n-K},\ldots,{token}_{n-1};\theta\right)$ (12)
@@ 5 Experiments
5.1 Experiment Setup
The experiments for the triage and consultation systems, including datasets, baselines, evaluation metrics, hyper-parameters, and environment settings, are outlined in detail in the Appendix C.
5.2 Supervised Triage System Experiment
5.2.1 Overall Performance
For the supervised learning classification method, we employ different baselines including machine learning methods, neural network methods, and pre-trained models. The results of different-grained datasets are shown in Table 3
Table 3: Experimental Results on Different Granularity Datasets
@@ 5.2.1 Overall Performance
Neural Network TextCNN 77.07% 54.36% 56.83% 52.10% 76.38% 61.29% 66.92% 56.54%
RCNN 71.56% 51.40% 53.76% 49.24% 73.84% 59.86% 61.10% 58.67%
@@
Ablation study results are presented in Appendix D.1
5.2.2 Results of Domain Information Injection
To investigate the differences before and after domain-specific re-training of different pre-trained models, we designed comparative experiments to evaluate model performance under different conditions. We used the original BERT and MacBERT models, as well as the proposed models utilizing these pre-trained models, comparing their performance before and after re-training. The results are shown in Figure 5
@@ 5.2.2 Results of Domain Information Injection
With the same pre-trained models and classification network structures, re-training the language model can improve the accuracy of the classification model. This is because the re-training data is related to Chinese medical information, allowing the language model to learn the intrinsic knowledge of medical texts, resulting in better performance on tasks involving medical texts.
@@
5.2.3 Parameters Study
We also investigated the impact of different numbers of dendritic neural network layers and LSTM layers on classification results to select the optimal algorithm model structure.
@@ 5.2.3 Parameters Study
According to the results in Figure 6, the dendritic neural network achieves the best classification performance with 3 layers, and the LSTM network achieves optimal performance with 2 layers.
5.3 Prompting Triage System Experiment
@@
Table 4: Prompt Learning Classification Results on Different Granularity Datasets
@@
Table 4 shows that the BERT model re-trained with collected data performs better in prompt learning-based classification, surpassing supervised learning on small datasets.
However, prompt learning has drawbacks: it requires more training time due to predicting multiple token positions, and its performance can be slightly worse than supervised learning. This is due to the need to correctly predict all token positions and handle varying label lengths with padding symbols, which can interfere with pre-trained knowledge and reduce accuracy. Additionally, prompt learning is less effective with large datasets.
5.4 Question Answer System Experiment
5.4.1 Overall Performance

### v2

# Building a Chinese Medical Dialogue System: Integrating Large-scale Corpora and Novel Models

@@ Abstract
The global COVID-19 pandemic underscored major deficiencies in traditional healthcare systems, hastening the advancement of online medical services, especially in medical triage and consultation. However, existing studies face two main challenges. First, the scarcity of large-scale, publicly available, domain-specific medical datasets due to privacy concerns, with current datasets being small and limited to a few diseases, limiting the effectiveness of triage methods based on Pre-trained Language Models (PLMs). Second, existing methods lack medical knowledge and struggle to accurately understand professional terms and expressions in patient-doctor consultations. To overcome these obstacles, we construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC) , thereby addressing the data shortage in this field. Moreover, we further propose a novel triage system that combines BERT-based supervised learning with prompt learning, as well as a GPT-based medical consultation model. To enhance domain knowledge acquisition, we pre-trained PLMs using our self-constructed background corpus. Experimental results on the LCMDC demonstrate the efficacy of our proposed systems. Our corpora and codes are available at: Corpora and Codes.
Index Terms:
Pre-trained Language Models, Medical Triage, Medical Consultation, Domain Knowledge Acquisition
I Introduction
@@
The outbreak and spread of COVID-19 have placed immense pressure on global healthcare systems, exposing numerous shortcomings in traditional healthcare services [1]. In response to the urgent need for pandemic control, online medical consultations and telemedicine have become crucial for maintaining healthcare services [2]. Between July and December 2022, 58.5% of U.S. adults used the Internet to seek health or medical information, 41.5% communicated with a doctor or doctor’s office, and 46.1% checked medical test results[3]. According to the China Internet Network Information Center (CNNIC), as of December 2023, the number of Internet medical users in China reached 414 million. This trend reflects the growing reliance on digital platforms for healthcare.
@@
The rise of online healthcare platforms has made medical triage and intelligent Q&A key research areas, helping reduce the burden on healthcare systems, providing remote medical advice, and reducing unnecessary hospital visits. These technologies enhance resource efficiency, minimize cross-infection risks, and improve access to healthcare services.
@@
However, most existing studies face two main challenges. Firstly, due to data privacy concerns, obtaining complete large-scale public medical triage and consultation datasets is difficult [4]. Although some studies have collected corpora for automatic diagnosis [5, 6, 7] and dialogue generation [8], these datasets are limited to a few specific diseases and are small in size [9], making them inadequate for training intelligent medical diagnosis models based on pre-trained language models. Secondly, patient-doctor consultations usually utilize distinct expressions and professional terms, making it difficult for existing methods that lack domain-specific knowledge to comprehend medically pertinent information [10].
To address the limitations, we have built a Chinese medical dialogue system, including 1) the medical corpora, 2) the medical triage system, and 3) the medical consultation model.
@@
The Large-scale Chinese Medical Dialogue Corpora (LCMDC) extracts patient-doctor conversations from the "Quick Doctor" website  [^1] https://www.120ask.com/. Our work addresses two key needs in online healthcare: task-oriented medical triage and open-ended consultation systems. We developed a coarse-grained triage dataset with over 430k patient-doctor dialogues across 14 departments and a fine-grained diagnosis dataset with 200k entries covering 120 diseases. For the consultation, we compiled over 470k question-answer pairs. Additionally, we collected disease information datasets to enrich Pre-trained Language Models (PLMs) with medical knowledge for pre-training.
@@
Nowadays, natural language processing techniques are increasingly effective in intelligent medical triage and dialogue generation tasks[11, 12]. Specifically, neural network-based text classification algorithms with attention mechanisms have been successfully applied to medical triage and diagnosis[5, 6, 7]. However, existing studies often concentrate on a limited number of diseases and struggle with imbalanced label distributions in many-class medical diagnoses [9]. Although some pre-trained language models, such as BERT and ERNIE, have been employed in automatic triage systems [13, 14], they lack pre-training on medical data, hindering their ability to grasp domain-specific knowledge [10]. To overcome these limitations, we propose a novel triage system that combines PLM-based supervised learning with prompt learning methods. We initially pre-train BERT using a self-built medical background corpus, then employ medical BERT, BiLSTM, and Dendrite Network [15] to represent the input questions for classification in our Triage datasets. Additionally, we incorporate prompt learning to mitigate the many-label imbalance issue, enhancing prediction accuracy for rare diseases.
@@
In the medical consultation model, sequence-to-sequence models, including LSTM-based Seq2Seq methods [8] and transformer-based PLMs like BERT-GPT [9], have been widely used, and achieved promising results. However, traditional LSTM models, which generate output solely based on input sequences, struggle with insufficient information utilization due to their gating mechanisms. Similarly, conventional PLMs often fail to effectively comprehend medical information within patient-doctor conversations. Additionally, these methods typically lack supplementary information for enhancement, resulting in poor performance in generating patient-doctor dialogues. To address these challenges, we re-train GPT-2 using our collected medical background corpus to internalize domain-specific knowledge. We also construct a domain knowledge graph and extract relevant content about each user’s question as supplementary input information to enhance the model. Finally, we train our model using our Medical Consultation Dataset.
@@
  • Developed a medical consultation model including knowledge incorporation and input supplementation, improving performance across multiple evaluation metrics.
II Related Work
The development of AI-driven medical dialogue systems benefits from advancements in biomedical research and feature engineering techniques. Studies on bioinformatics and medicine [16, 17, 18, 19, 20, 21] provide valuable domain knowledge that can enhance medical AI applications. Additionally, effective feature selection and transformation techniques are crucial for optimizing model performance. Feature selection methods [22, 23, 24, 25] help identify the most informative clinical attributes, while feature transformation approaches [26, 27, 28, 29, 30] improve data representation, enabling more accurate and robust medical triage and consulting systems. Furthermore, large language models (LLMs) have significantly advanced research in medical AI by improving the understanding and processing of complex biomedical texts [31, 32, 33]. Pre-trained models with domain-specific fine-tuning enable more effective medical dialogue generation, question answering, and clinical decision support.
II-A Dataset
CBLUE [34] is a comprehensive Chinese biomedical language understanding benchmark designed to advance research in biomedical NLP for the Chinese language. MedDG [8] is a large-scale dataset aimed at developing conversational agents for primary clinical advice, especially valuable during the COVID-19 pandemic. MedDialog [9] is a comprehensive dataset for training and evaluating medical dialogue systems, enhancing remote healthcare services. MedQuAD [35] is a significant resource for developing AI models in medical question answering, providing a robust foundation for understanding and responding to medical inquiries. There are some other datasets like event recognition [36].
II-B Triage System
IntelTriage [37] is an intelligent triage system for Emergency Departments in Greece, dynamically prioritizing patients and monitoring their vital signs using wearable biosensors. The Babylon AI-powered Triage and Diagnostic System [38] demonstrated diagnostic accuracy comparable to human doctors and provided safer triage recommendations. MESTRIMAN [39] is an expert system designed to support medical assistance during catastrophes by coordinating resources and enhancing disaster management. The SADE model [10] addresses challenges in medical triage and diagnosis by using dual encoders to represent patient consultations and doctor diagnoses.
II-C Question Answer System
This research [40] automates the generation of qualified medical answers using an end-to-end RNN model trained on data from online health services. Another system [41] integrates medical knowledge, knowledge graphs, and QA systems, using crawler technology and rule-based algorithms for classification and querying. An end-to-end character-level multi-scale CNN framework [42] is introduced for Chinese medical question-answer matching. Med-PaLM 2 [43] leverages large language model improvements and medical domain fine-tuning to enhance performance on medical QA datasets.
III Chinese Medical Corpus
@@
III-A Data Collection
To make the model more adaptable to conversational and ambiguous medical descriptions, we claw the original data from the online medical consultation platform named "Quick Doctor". Here are several reasons:
@@ III-A Data Collection
We collect millions of entries containing doctor responses and conduct statistical analysis and visualization based on different labels to better understand the original data. Our analysis, illustrated in Figure 1, explores the gender and age of users. The results show that female users significantly outnumber male users. Predominantly, young individuals aged 10-30, who are more likely to incorporate the Internet into their daily routines, made the inquiries.
III-B Intelligent Triage Dataset
@@
We organize the original corpus into two datasets after data preprocessing: 1) a Coarse-grained Triage dataset with primary department labels for higher accuracy in quickly identifying potential conditions and directing patients to the appropriate department, and 2) a Fine-grained Diagnosis dataset with tertiary department labels for more precise guidance, thereby helping users access detailed medical knowledge and improving query efficiency. The Coarse-grained dataset has over 430k samples with 14 categories, while the Fine-grained dataset has over 190k samples with 120 categories, as shown in Table I. Based on these datasets, we further propose an intelligent triage algorithm based on supervised learning and prompt-based learning and then evaluate its effectiveness.
TABLE I: Statistic of Intelligent Triage Datasets
@@
III-C Medical Consultation Dataset
@@
TABLE II: Statistic of Medical Consulting Dataset
@@
Specifically, each sample in this dataset comprises a patient consultation and a doctor’s response, serving respectively as the model’s input and prediction target. We conducted data cleaning to remove entries that lacked either questions or answers and eliminated entries where the questions and responses were shorter than 10 characters, thereby enriching the training data with more substantive content. After data preprocessing, the Medical Consultation Dataset contains a total of 472,418 samples, with an average question length of 57.86 characters and an average response length of 92.57 characters. The statistical results are presented in Table II.
IV Intelligent Triage System
@@ IV-A Problem Definition
IV-B Supervised Learning Triage System
@@
IV-C Prompt Learning Triage System
@@
V Medical Consultation System
@@ V-A Problem Definition
V-B Framework Overview
Based on the GPT-2 model, we re-train the model using the collected medical encyclopedia text data to internalize knowledge, making it more suitable for tasks in the medical domain. Next, we construct a medical knowledge graph and use a search-and-match approach to find relevant knowledge to each question, providing external information as supplementary input text. Then, we train the model using dialogue data.
@@ V-B Framework Overview
V-C Knowledge Injection Answer Generation
@@
$\scriptsize L\left(Sentence\right)=\sum\limits_{n=K}^{N}p\left({token}_{n}\,\middle|\,{token}_{n-K},\ldots,{token}_{n-1};\theta\right)$ (12)
@@ VI Experiments
VI-A Experiment Setup
VI-A1 Datasets
We used the datasets shown in Section III, including the intelligent triage and consultant datasets.
VI-A2 Baselines
We employ the mechanism learning baselines, neural network baselines, and pre-trained language models, shown in Table III and Table V.
VI-A3 Evaluation Metrics
For the triage system, we utilize accuracy, precision, recall, and F-1 score. For the consultation system, we employ n-gram overlap, distance-based, vector similarity, and sentence semantic similarity measurements.
VI-A4 Hyper Parameters
Continuous pre-training: Max length 128, mask 15% tokens in MLM. Supervised training: Embedding dim 768, LSTM hidden size 1024, 50 epochs. PTM lr 5e-5, others 2e-4. Prompt learning: 20 epochs, lr 2e-5. Consultation system: Base model GPT-2, token embedding dim 768, vocab length 50257. Pre-train 10 epochs with background knowledge, fine-tune 20 epochs, lr 2.6e-5.
VI-A5 Experimental Environment
VI-B Supervised Triage System Experiment
VI-B1 Overall Performance
For the supervised learning classification method, we employ different baselines including machine learning methods, neural network methods, and pre-trained models. The results of different-grained datasets are shown in Table III.
TABLE III: Experimental Results on Different Granularity Datasets
@@ VI-B1 Overall Performance
Neural Network TextCNN [44] 77.07% 54.36% 56.83% 52.10% 76.38% 61.29% 66.92% 56.54%
RCNN [45] 71.56% 51.40% 53.76% 49.24% 73.84% 59.86% 61.10% 58.67%
@@
VI-B2 Ablation study
(a) Coarse-grained
(b) Fine-grained
Figure 5: Ablation Results of Supervised Classification Method.
VI-B3 Results of Domain Information Injection
To investigate the differences before and after domain-specific re-training of different pre-trained models, we designed comparative experiments to evaluate model performance under different conditions. We used the original BERT and MacBERT models, as well as the proposed models utilizing these pre-trained models, comparing their performance before and after re-training. The results are shown in Figure 6.
@@ VI-B3 Results of Domain Information Injection
With the same pre-trained models and classification network structures, re-training the language model can improve the classification accuracy. This is because the re-training data is related to Chinese medical information, allowing the language model to learn the intrinsic knowledge of medical texts, resulting in better performance of medical texts.
@@
VI-B4 Parameters Study
We also investigated the impact of different numbers of layers of dendritic neural network and LSTM layers on classification results to select the optimal algorithm model structure.
@@ VI-B4 Parameters Study
According to Figure 7, the dendritic neural network achieves the best classification performance with 3 layers, and the LSTM network achieves optimal performance with 2 layers.
VI-C Prompting Triage System Experiment
@@
TABLE IV: Prompt Learning Classification Results
@@
Table IV shows that the BERT model re-trained with collected data performs better in prompt learning-based classification, surpassing supervised learning on small datasets.
However, prompt learning has drawbacks: it requires more time due to predicting multiple tokens, and its performance can be slightly worse than supervised learning. This is due to the need to correctly predict all token positions and handle varying label lengths with padding symbols, which can interfere with pre-trained knowledge and reduce accuracy. Additionally, prompt learning is less effective with large datasets.
VI-D Medical Consultation System Experiment
VI-D1 Overall Performance